OpenStudy (anonymous):

present value of future amount $400,000 at 11% compounded semi annually for 25 years

OpenStudy (anonymous):

can you please reword the question?

OpenStudy (anonymous):

Find the present value of the following future amount. $400,000 at 11% compounded semi annually for 25 years. Thank you

OpenStudy (wolf1728):

Basically, you want to know the amount of principal that grwos to 400,000 at 11% for 25 years?

OpenStudy (anonymous):

Yes

OpenStudy (wolf1728):

The formula is here: http://1728.org/compint3.htm but I can solve it for you. Principal = Total /(1 + Rate/n) ^ n*years Principal = 400,000 / 1+ (.11/2)^2*25 Principal = 400,000 / 1.055)^50 Principal = 400,000 / 14.5419612045 Principal = 27,506.61

OpenStudy (anonymous):

Thank you!

OpenStudy (wolf1728):

u r welcome :-)